# SemiETPicker: Fast and Label-Efficient Particle Picking for CryoET Tomography Using Semi-Supervised Learning

### v1

@@ 1 Introduction
Cryogenic Electron Tomography (CryoET) enables the visualization of macromolecular complexes in near-native conformations at sub-nanometer resolution [1]. In a typical CryoET pipeline, frozen hydrated samples are imaged by tilting them incrementally under an electron beam to acquire a tilt-series of 2D projections. These projections are computationally reconstructed into a 3D density map known as a tomogram. Downstream structural analysis critically relies on particle picking, i.e., the localization and classification of sub-cellular components within the tomogram. Existing approaches [2, 3, 4, 5] typically rely on supervised learning or template matching [6], both of which require extensive manual annotations. To alleviate annotation costs, recent efforts have explored few-shot learning [7] and contrastive learning [8]. However, the wealth of unlabeled CryoET tomograms remains under-utilized.
Semi-Supervised Learning (SSL) leverages a small labeled dataset and a large unlabeled dataset to improve model performance. Landmark methods in image classification, such as Mean Teacher [9] and FixMatch [10], demonstrate that enforcing prediction consistency between weak and strong augmentations using high-confidence pseudo labels can yield substantial gains. In particular, FixMatch simplifies the SSL pipeline by employing the teacher model to generate pseudo labels from weakly augmented images, which are then used to supervise the student model trained on strongly augmented versions. Only pseudo labels with high confidence are retained, making this a simple and effective strategy in modern SSL.
Extending SSL to object detection introduces additional challenges due to class imbalance and the dual tasks of classification and localization [11, 12, 13]. While particle picking bears similarities to natural image object detection, it presents unique difficulties: (1) the low signal-to-noise ratio in tomograms renders conventional deep architectures suboptimal; (2) confidence-based pseudo-label filtering alone fails to ensure quality supervision; and (3) strong augmentations effective for CryoET images, characterized by small, densely packed particles, are underexplored.
@@ 2.1 Problem Description
Currently, a large number of CryoET tomograms have been generated. However, each tomogram contains hundreds of protein particles. Labeling such densely distributed and small particles is very time-consuming, so only a small fraction are annotated [14]. We denote the labeled dataset as $\mathcal{D}^{L}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{l}$, where $l\ll u$, and evaluation is performed on separate tomograms. The unlabeled dataset is denoted as $\mathcal{D}^{U}=\{(\mathbf{x}^{u}_{i})\}_{i=1}^{u}$, where $\mathbf{x}_{i}\in\mathbb{R}^{D\times H\times W}$ are tomograms and $y$ are labels. In the semi-supervised setting, we train our model jointly on $\mathcal{D}^{L}$ and $\mathcal{D}^{U}$.
@@ 2.2 An End-to-End Object Detector for Particle-Picking
Here we reformulate the particle picking problem as a keypoint prediction problem inspired by [15, 16]. Let $I\in\mathcal{R}^{D\times H\times W}$ be an input tomogram of depth $D$, height $H$, and width $W$. We aim to predict a heatmap $\hat{Y}\in[0,1]^{C\times\frac{D}{R}\times\frac{H}{R}\times\frac{W}{R}}$, where $R$ is the output stride and $C$ is the number of particle types. We experimentally find $R=2$ achieves a good trade-off between performance and speed. We design a asymmetric UNet accordingly to match the output size [17].
@@
where $\sigma_{c}$ is the standard deviation, set to half the typical radius of particles of class $c$. MSE loss is a common choice for heatmap regression [18, 19]. Due to the predominance of background voxels in CryoET volumes, we propose a reweighted MSE loss to address the severe class imbalance:
@@
where $\lambda=4$ is a weighting factor. For clarity, we omit the subscripts czyxczyx in the equation. Empirically, we find that using a larger $\lambda$ improves performance, likely due to its role in emphasizing hard negative examples [20].
@@ 2.3 Teacher-Student Co-training and Multi-View Pseudo Labeling
The second stage is a teacher-student co-training stage. We duplicate the weights $\theta$ to initialize both the teacher model $\theta_{t}$ and the student model $\theta_{s}$. The student model is trained using both labeled data $\mathcal{D}^{L}$ and unlabeled data $\mathcal{D}^{U}$. Meanwhile, the teacher model is updated via Exponential Moving Average (EMA) of the student weights: $\theta_{t}\leftarrow(1-\alpha)\theta_{t}+\alpha\theta_{s},$ where $\alpha$ is a smoothing coefficient. As demonstrated in prior work [9], the EMA-updated teacher consistently outperforms the student model in prediction quality and is thus used to generate supervision signals.
@@ 2.4 CryoET-specific DropBlock Augmentation
In the co-training process, the choice of strong augmentations $\mathcal{A}_{s}$ is pivotal. While strong augmentations such as RandAugment [21], MixUp [22], and CutOut [23] have proven effective in natural image domains, their direct application to CryoET is suboptimal due to the distinct characteristics of tomographic data.
@@
In natural images, foreground objects are often large and sparse. CutOut, which removes random rectangular patches, mimics occlusions—a realistic challenge in object detection. In contrast, CryoET features small, densely packed particles. Applying CutOut often removes entire particles, depriving the model of critical training information. To address this, we adapt DropBlock [24] for CryoET data. Originally proposed as a regularization technique for CNNs, DropBlock randomly removes contiguous regions of feature maps during training. In our adaptation, we first sample a binary mask from a Bernoulli distribution with the same shape as the image input. We then apply max pooling with a small kernel(size=3) and a stride equal to half the kernel size to expand the dropped regions. Finally, we threshold the result to produce a binary mask that selectively occludes parts of particles while preserving sufficient structural context for effective learning. In our experiments, integrating this tailored DropBlock variant into $\mathcal{A}_{s}$ significantly boosts performance in the semi-supervised setting.
@@ 3.1 Dataset
We use the CZII dataset [14] to evaluate our algorithm because it is the largest CryoET dataset and provides multi-class annotations. It includes center annotations for six particle categories, with 7 tomograms for training and 500 for testing. The significant imbalance stems from the high cost of annotating multiple particle types across many tomograms, which is infeasible for most biology labs [14]. To simulate this scenario, we randomly select 2 and 5 tomograms as the labeled training set $\mathcal{D}^{L}$, 50 tomograms as the unlabeled set $\mathcal{D}^{U}$, and another 50 tomograms as the test set $\mathcal{D}^{T}$. Each tomogram has dimensions $183\times 650\times 650$ with a voxel spacing of 10 Å. The six particle classes have radii ranging from 60 Å to 150 Å, representative of most particle sizes observed in practice, with typical radii shown in Table 2. The classes are apo-ferritin (AF), beta-amylase (BA), beta-galactosidase (BG), ribosome (Ribo), thyroglobulin (Thy), and virus-like particle (VLP).
@@ 3.2 Implementation Details
For training, we crop subtomograms of size $90\times 90\times 90$ as input. Our asymmetric U-Net [17] and ResNet-34 [25] backbones are adapted with 3D operators. Specifically, we replace the initial $7\times 7\times 7$ convolutional stem in ResNet with three successive $3\times 3\times 3$ convolutions, following [26]. Models are trained using the Adam optimizer [27] on two NVIDIA A30 GPUs. Intensity augmentations are applied using the MONAI framework [28]. During inference, we employ a sliding window strategy with 0.25 overlap, and apply test-time augmentation via three-axis flipping.
@@
DeepFindET [29, 30] 0.241 0.250 0.328 0.341
VNet [31] 0.507 0.518 0.564 0.574
Mean Teacher [9] 0.527 0.573 0.586 0.633
@@ 3.3 Comparison with Other Methods
We compare our method against a range of baselines. First, we evaluate it against Supervised, which is trained on $\mathcal{D}^{L}$ with our asymmetric UNet and reweighted MSE loss. This comparison demonstrates how effectively our semi-supervised learning pipeline leverages unlabeled data. To verify the validity of the Supervised, we also compare it with other supervised methods such as DeepFindET [29, 30] and VNet [31]. Notably, DeepFindET is specifically optimized for Cryo-ET particle picking. Additionally, we implement a classical semi-supervised learning method, Mean Teacher [9], as a baseline. To illustrate the upper bound of performance, we include a Supervised full baseline trained on both $\mathcal{D}^{L}$ and $\mathcal{D}^{U}$ with ground truth labels.
@@ 3.4 Ablation Studies and Efficiency Analysis
DeepFindET [29] 41s 0.328
VNet [31] 2.66s 0.564
FPN [32] 0.58s 0.547
@@
Another advantage of SemiETPicker is its efficiency. This efficiency stems from two design choices. First, our model outputs a heatmap, and the post-processing involves a simple max pooling operation, which can be efficiently accelerated by GPUs. As shown in Table 4, DeepFindET is significantly slower than the other three methods. This is because DeepFindET adopts a segmentation-based design, requiring Connected Component Labeling (CCL) [33] as a post-processing step to extract particle centers. This operation runs on the CPU and is not well-suited for parallel computation, leading to a substantial bottleneck in the pipeline. Second, our asymmetric UNet design offers a favorable speed–accuracy trade-off. In Table 4, VNet and FPN [32] are trained and evaluated with the same pipeline as our supervised model, differing only in the backbone. This isolates the effect of postprocessing. FPN runs faster without skip connections but performs much worse. VNet, which predicts at input resolution, matches our accuracy but is far slower.

### v2

@@ 1 Introduction
Cryogenic Electron Tomography (CryoET) enables the visualization of macromolecular complexes in near-native conformations at sub-nanometer resolution [doerr2017cryo]. In a typical CryoET pipeline, frozen hydrated samples are imaged by tilting them incrementally under an electron beam to acquire a tilt-series of 2D projections. These projections are computationally reconstructed into a 3D density map known as a tomogram. Downstream structural analysis critically relies on particle picking, i.e., the localization and classification of sub-cellular components within the tomogram. Existing approaches [wagner2019sphire, hao2022vp, gubins2020shrec, zhou2023machine] typically rely on supervised learning or template matching [frangakis2002identification], both of which require extensive manual annotations. To alleviate annotation costs, recent efforts have explored few-shot learning [zhou2021one] and contrastive learning [huang2022accurate]. However, the wealth of unlabeled CryoET tomograms remains under-utilized.
Semi-Supervised Learning (SSL) leverages a small labeled dataset and a large unlabeled dataset to improve model performance. Landmark methods in image classification, such as Mean Teacher [tarvainen2017mean] and FixMatch [sohn2020fixmatch], demonstrate that enforcing prediction consistency between weak and strong augmentations using high-confidence pseudo labels can yield substantial gains. In particular, FixMatch simplifies the SSL pipeline by employing the teacher model to generate pseudo labels from weakly augmented images, which are then used to supervise the student model trained on strongly augmented versions. Only pseudo labels with high confidence are retained, making this a simple and effective strategy in modern SSL.
Extending SSL to object detection introduces additional challenges due to class imbalance and the dual tasks of classification and localization [sohn2020simple, xu2021end, liu2021unbiased]. While particle picking bears similarities to natural image object detection, it presents unique difficulties: (1) the low signal-to-noise ratio in tomograms renders conventional deep architectures suboptimal; (2) confidence-based pseudo-label filtering alone fails to ensure quality supervision; and (3) strong augmentations effective for CryoET images, characterized by small, densely packed particles, are underexplored.
@@ 2.1 Problem Description
Currently, a large number of CryoET tomograms have been generated. However, each tomogram contains hundreds of protein particles. Labeling such densely distributed and small particles is very time-consuming, so only a small fraction are annotated [peck2024annotating]. We denote the labeled dataset as $\mathcal{D}^{L}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{l}$, where $l\ll u$, and evaluation is performed on separate tomograms. The unlabeled dataset is denoted as $\mathcal{D}^{U}=\{(\mathbf{x}^{u}_{i})\}_{i=1}^{u}$, where $\mathbf{x}_{i}\in\mathbb{R}^{D\times H\times W}$ are tomograms and $y$ are labels. In the semi-supervised setting, we train our model jointly on $\mathcal{D}^{L}$ and $\mathcal{D}^{U}$.
@@ 2.2 An End-to-End Object Detector for Particle-Picking
Here we reformulate the particle picking problem as a keypoint prediction problem inspired by [law2018cornernet, zhou2019objects]. Let $I\in\mathcal{R}^{D\times H\times W}$ be an input tomogram of depth $D$, height $H$, and width $W$. We aim to predict a heatmap $\hat{Y}\in[0,1]^{C\times\frac{D}{R}\times\frac{H}{R}\times\frac{W}{R}}$, where $R$ is the output stride and $C$ is the number of particle types. We experimentally find $R=2$ achieves a good trade-off between performance and speed. We design a asymmetric UNet accordingly to match the output size [ronneberger2015u].
@@
where $\sigma_{c}$ is the standard deviation, set to half the typical radius of particles of class $c$. MSE loss is a common choice for heatmap regression [luo2021rethinking, uchida20252]. Due to the predominance of background voxels in CryoET volumes, we propose a reweighted MSE loss to address the severe class imbalance:
@@
where $\lambda=4$ is a weighting factor. For clarity, we omit the subscripts czyxczyx in the equation. Empirically, we find that using a larger $\lambda$ improves performance, likely due to its role in emphasizing hard negative examples [shrivastava2016training].
@@ 2.3 Teacher-Student Co-training and Multi-View Pseudo Labeling
The second stage is a teacher-student co-training stage. We duplicate the weights $\theta$ to initialize both the teacher model $\theta_{t}$ and the student model $\theta_{s}$. The student model is trained using both labeled data $\mathcal{D}^{L}$ and unlabeled data $\mathcal{D}^{U}$. Meanwhile, the teacher model is updated via Exponential Moving Average (EMA) of the student weights: $\theta_{t}\leftarrow(1-\alpha)\theta_{t}+\alpha\theta_{s},$ where $\alpha$ is a smoothing coefficient. As demonstrated in prior work [tarvainen2017mean], the EMA-updated teacher consistently outperforms the student model in prediction quality and is thus used to generate supervision signals.
@@ 2.4 CryoET-specific DropBlock Augmentation
In the co-training process, the choice of strong augmentations $\mathcal{A}_{s}$ is pivotal. While strong augmentations such as RandAugment [cubuk2020randaugment], MixUp [zhang2017mixup], and CutOut [devries2017improved] have proven effective in natural image domains, their direct application to CryoET is suboptimal due to the distinct characteristics of tomographic data.
@@
In natural images, foreground objects are often large and sparse. CutOut, which removes random rectangular patches, mimics occlusions—a realistic challenge in object detection. In contrast, CryoET features small, densely packed particles. Applying CutOut often removes entire particles, depriving the model of critical training information. To address this, we adapt DropBlock [ghiasi2018dropblock] for CryoET data. Originally proposed as a regularization technique for CNNs, DropBlock randomly removes contiguous regions of feature maps during training. In our adaptation, we first sample a binary mask from a Bernoulli distribution with the same shape as the image input. We then apply max pooling with a small kernel(size=3) and a stride equal to half the kernel size to expand the dropped regions. Finally, we threshold the result to produce a binary mask that selectively occludes parts of particles while preserving sufficient structural context for effective learning. In our experiments, integrating this tailored DropBlock variant into $\mathcal{A}_{s}$ significantly boosts performance in the semi-supervised setting.
@@ 3.1 Dataset
We use the CZII dataset [peck2024annotating] to evaluate our algorithm because it is the largest CryoET dataset and provides multi-class annotations. It includes center annotations for six particle categories, with 7 tomograms for training and 500 for testing. The significant imbalance stems from the high cost of annotating multiple particle types across many tomograms, which is infeasible for most biology labs [peck2024annotating]. To simulate this scenario, we randomly select 2 and 5 tomograms as the labeled training set $\mathcal{D}^{L}$, 50 tomograms as the unlabeled set $\mathcal{D}^{U}$, and another 50 tomograms as the test set $\mathcal{D}^{T}$. Each tomogram has dimensions $183\times 650\times 650$ with a voxel spacing of 10 Å. The six particle classes have radii ranging from 60 Å to 150 Å, representative of most particle sizes observed in practice, with typical radii shown in Table 2. The classes are apo-ferritin (AF), beta-amylase (BA), beta-galactosidase (BG), ribosome (Ribo), thyroglobulin (Thy), and virus-like particle (VLP).
@@ 3.2 Implementation Details
For training, we crop subtomograms of size $90\times 90\times 90$ as input. Our asymmetric U-Net [ronneberger2015u] and ResNet-34 [he2016deep] backbones are adapted with 3D operators. Specifically, we replace the initial $7\times 7\times 7$ convolutional stem in ResNet with three successive $3\times 3\times 3$ convolutions, following [he2019bag]. Models are trained using the Adam optimizer [kingma2014adam] on two NVIDIA A30 GPUs. Intensity augmentations are applied using the MONAI framework [cardoso2022monai]. During inference, we employ a sliding window strategy with 0.25 overlap, and apply test-time augmentation via three-axis flipping.
@@
DeepFindET [moebel2021deep, harrington2024open] 0.241 0.250 0.328 0.341
VNet [milletari2016v] 0.507 0.518 0.564 0.574
Mean Teacher [tarvainen2017mean] 0.527 0.573 0.586 0.633
@@ 3.3 Comparison with Other Methods
We compare our method against a range of baselines. First, we evaluate it against Supervised, which is trained on $\mathcal{D}^{L}$ with our asymmetric UNet and reweighted MSE loss. This comparison demonstrates how effectively our semi-supervised learning pipeline leverages unlabeled data. To verify the validity of the Supervised, we also compare it with other supervised methods such as DeepFindET [moebel2021deep, harrington2024open] and VNet [milletari2016v]. Notably, DeepFindET is specifically optimized for Cryo-ET particle picking. Additionally, we implement a classical semi-supervised learning method, Mean Teacher [tarvainen2017mean], as a baseline. To illustrate the upper bound of performance, we include a Supervised full baseline trained on both $\mathcal{D}^{L}$ and $\mathcal{D}^{U}$ with ground truth labels.
@@ 3.4 Ablation Studies and Efficiency Analysis
DeepFindET [moebel2021deep] 41s 0.328
VNet [milletari2016v] 2.66s 0.564
FPN [lin2017feature] 0.58s 0.547
@@
Another advantage of SemiETPicker is its efficiency. This efficiency stems from two design choices. First, our model outputs a heatmap, and the post-processing involves a simple max pooling operation, which can be efficiently accelerated by GPUs. As shown in Table 4, DeepFindET is significantly slower than the other three methods. This is because DeepFindET adopts a segmentation-based design, requiring Connected Component Labeling (CCL) [horn1986robot] as a post-processing step to extract particle centers. This operation runs on the CPU and is not well-suited for parallel computation, leading to a substantial bottleneck in the pipeline. Second, our asymmetric UNet design offers a favorable speed–accuracy trade-off. In Table 4, VNet and FPN [lin2017feature] are trained and evaluated with the same pipeline as our supervised model, differing only in the backbone. This isolates the effect of postprocessing. FPN runs faster without skip connections but performs much worse. VNet, which predicts at input resolution, matches our accuracy but is far slower.